### v2

@@ 4.2 Deriving the strong second-order sufficient condition
$\begin{array}[]{ll}H^{\top}\Theta HP_{k}\tilde{\bm{\eta}}+\sum\limits_{i=k+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}\tilde{\boldsymbol{p}}^{j_{i}}\in\ker(P_{k}^{\top})={\sf span}(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}),\end{array}$
which contradicts (4.35). Therefore, (4.34) holds, and it is easy to see that $T_{k}$ is nonsingular.
In the following, we separate the proof of (4.30) into three cases:
@@
Let $k:=\iota-1$ in (4.32). According to (2.4) and Part (6e) of Lemma 3.2, one can fix all $p^{j_{i}}=0$ and all ${\boldsymbol{p}}^{j_{i}}={\boldsymbol{0}}$ in (4.32) for all $i\neq\iota$. By doing this one can obtain that
$T_{\iota-1}{\bm{\eta}}\neq-P_{\iota-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}p,\quad\forall\,p\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{\iota}}P_{\iota-1}{\bm{\eta}}),\quad\forall\,{\bm{\eta}}\in{\mathds{R}}^{\ell-\iota+1}\setminus\{{\boldsymbol{0}}\}.$ (4.36)
From (4.25) one has ${\mathcal{G}}_{j_{\iota}}=((\dot{\mu}^{*})^{j_{\iota}};-(\bar{\bm{\mu}}^{*})^{j_{\iota}})^{\top}{\mathcal{J}}g^{j_{\iota}}({\boldsymbol{x}}^{*})H$. Then, by defining the vector ${\bm{\eta}}:=T_{\iota-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{\iota-1}]^{\top}$ one can get
$\langle(\dot{\mu}^{*})^{j_{\iota}};-(\bar{\bm{\mu}}^{*})^{j_{\iota}}),{\mathcal{J}}g^{j_{\iota}}({\boldsymbol{x}}^{*})HP_{{\iota}-1}{\bm{\eta}}\rangle={\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}{\bm{\eta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\leq 0.$ (4.37)
It can be easily observed from (2.4) that $p:=-1\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}{\bm{\eta}})$. Meanwhile, one has $T_{{\iota}-1}{\bm{\eta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}=-P_{{\iota}-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}p$. However, this contradicts (4.36), so that (4.37) fails. Therefore, (4.30) holds for all $i\in\{1,\ldots,|J_{4}|\}$.
@@
Let $k:=\iota-1$ in (4.32). According to (2.4) and Part (6e) of Lemma 3.2, one can fix all $p^{j_{i}}=0$ and all ${\boldsymbol{p}}^{j_{i}}={\boldsymbol{0}}$ in (4.32) for all $i\neq\iota$. By doing this one can obtain that
$T_{\iota-1}{\bm{\eta}}\neq-P_{\iota-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}p,\quad\forall\,p\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{\iota}}P_{\iota-1}{\bm{\eta}}),\quad\forall\,{\bm{\eta}}\in{\mathds{R}}^{\ell-\iota+1}\setminus\{{\boldsymbol{0}}\}.$ (4.38)
From (4.25) one has ${\mathcal{G}}_{j_{\iota}}=(\dot{g}^{j_{\iota}}({\boldsymbol{x}}^{*});-\bar{g}^{j_{\iota}}({\boldsymbol{x}}^{*}))^{\top}{\mathcal{J}}g^{j_{\iota}}({\boldsymbol{x}}^{*})H$. Then, by letting ${\bm{\eta}}:=T_{\iota-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{\iota-1}]^{\top}$ one can get
$(\dot{g}^{j_{\iota}}({\boldsymbol{x}}^{*});-\bar{g}^{j_{\iota}}({\boldsymbol{x}}^{*}))^{\top}{\mathcal{J}}g^{j_{\iota}}({\boldsymbol{x}}^{*})HP_{{\iota}-1}{\bm{\eta}}={\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}{\bm{\eta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\leq 0.$ (4.39)
Thus, it can be easily observed from (2.4) that $p:=-1\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}{\bm{\eta}})$. Moreover, by the definition of ${\bm{\eta}}$ we have $T_{{\iota}-1}{\bm{\eta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}=P_{{\iota}-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}=-P_{{\iota}-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}p$. However, this contradicts (4.38), so that (4.39) fails. Therefore, (4.30) holds for all $i\in\{|J_{4}|+1,\ldots,|J_{4,5}|\}$.
@@
$[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\not\succ O.$ (4.40)
Let $k:=\iota-1$ in (4.34). According to Part (6e) of Lemma 3.2 (or (2.4) if $r_{j_{\iota}=0}$), one can fix ${\boldsymbol{p}}^{j_{i}}={\boldsymbol{0}}$ in (4.34) for all $i\neq\iota$. By doing this and using the fact that $\tilde{k}=k+\sum_{t=|J_{4,5}|+1}^{k}r_{j_{t}}$ one can obtain that
$T_{\iota-1}{\bm{\eta}}\neq-P_{\iota-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}{\boldsymbol{p}},\quad\forall\,{\boldsymbol{p}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{\iota}}P_{\iota-1}{\bm{\eta}}),\quad\forall\,{\bm{\eta}}\in{\mathds{R}}^{\ell-\tilde{k}}\setminus\{{\boldsymbol{0}}\}.$ (4.41)
Note that $\ker P_{\iota-1}^{\top}={\sf span}\big{(}\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,\iota-1}\big{)}$. Moreover, recall from (4.27) and (4.28) that each $P_{i}$ has full column rank. Suppose that $[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}$ is singular, i.e., there exists a nonzero vector $\hat{\bm{\theta}}\in{\mathds{R}}^{1+r_{j_{\iota}}}$ such that $[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\hat{\bm{\theta}}={\boldsymbol{0}}.$ Then, on the one hand, if $\hat{\bm{\eta}}:=T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\hat{\bm{\theta}}={\boldsymbol{0}}$, one has $P^{\top}_{{\iota}-1}{\mathcal{G}}_{j_{\iota}}^{\top}\hat{\bm{\theta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\hat{\bm{\theta}}={\boldsymbol{0}}$, so that
@@
but this contradicts (4.20) of Corollary 4.2. On the other hand, if $\hat{\bm{\eta}}\neq{\boldsymbol{0}}$, one has from the fact ${\mathcal{G}}_{j_{\iota}}P_{\iota-1}\hat{\bm{\eta}}={\boldsymbol{0}}$ and Part (6a) of Lemma 3.2 (or (2.4) if $r_{j}=0$) that $-\hat{\bm{\theta}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{\iota}}P_{\iota-1}\hat{\bm{\eta}})$. Moreover, one can see that $T_{\iota-1}\hat{\bm{\eta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\hat{\bm{\theta}}=-P_{\iota-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}(-\hat{\bm{\theta}})$, which contradicts (4.41). Therefore, $[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}$ is nonsingular. Then, one can define the matrix $M:=([{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top})^{-1}$.
@@
For a vector ${\boldsymbol{q}}\in{\mathds{R}}^{1+r_{j_{\iota}}}$, one can define ${\boldsymbol{p}}:=-M{\boldsymbol{q}}$ and ${\bm{\eta}}:=-T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}{\boldsymbol{p}}$. Then one has ${\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}{\bm{\eta}}={\boldsymbol{q}}$ and, consequently, $T_{{\iota}-1}{\bm{\eta}}=-[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}{\boldsymbol{p}}=-P_{{\iota}-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}{\boldsymbol{p}}$. Moreover, since (4.41) holds, one obtains that
$-M{\boldsymbol{q}}\notin{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})({\boldsymbol{q}}),\quad\forall{\boldsymbol{q}}\in{{\mathds{R}}^{1+r_{{j}_{\iota}}}}.$ (4.42)
@@
Next, we show that (4.40) can not be true. We first consider the case that $r_{j_{\iota}}=0$. If (4.40) holds in this case, by setting $q=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}<0$, one has $-M^{-1}q=-1<0$. Then, by (2.4) we know that $-Mq\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}(0,0)(q)$, which contradicts (4.42). Consequently, one has $[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}>0$ in this case. In the following, we deal with the case that $r_{j_{\iota}}\geq 1$. Firstly, we shall prove that
@@
$\langle\bm{\theta},M^{-1}\bm{\theta}\rangle>0,\quad\forall\,\bm{\theta}\in{\sf bdry}{\mathcal{Q}}_{{j_{\iota}}}\setminus\{{\boldsymbol{0}}\}.$ (4.43)
@@
Suppose on the contrary that there exists a vector $\hat{\bm{\theta}}\in{\sf bdry}{\mathcal{Q}}_{{j_{\iota}}}\setminus\{{\boldsymbol{0}}\}$ such that $\langle\hat{\bm{\theta}},M^{-1}\hat{\bm{\theta}}\rangle\leq 0.$ By taking $\hat{\boldsymbol{q}}:=M^{-1}\bf\hat{\bm{\theta}}$, one has $\langle-\hat{\bm{\theta}},\hat{\boldsymbol{q}}\rangle\geq 0$. Then, by Part (6c) of Lemma 3.2 we know that $-M\hat{\boldsymbol{q}}=-\hat{\bm{\theta}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})(\hat{\boldsymbol{q}})$, which contradicts (4.42). Therefore, (4.43) holds. Secondly, we shall show that the following holds:
$\langle\bm{\theta},M^{-1}\bm{\theta}\rangle>0,\quad\forall\,\bm{\theta}\in{\sf int}{\mathcal{Q}}_{j_{\iota}}.$ (4.44)
@@
For the purpose of contradiction, assume that (4.44) does not hold. Since the matrix $M^{-1}$ is nonsingular and symmetric, we know from (4.43) and Lemma 2.1 that there exists a nonzero vector $\tilde{\boldsymbol{p}}\in-{\sf int}{\mathcal{Q}}_{j_{\iota}}$ such that $\tilde{\boldsymbol{q}}:=-M^{-1}\tilde{\boldsymbol{p}}\in-{\mathcal{Q}}_{j_{\iota}}$. From Part (6b) of Lemma 3.2 one can see that $-M\tilde{\boldsymbol{q}}=\tilde{\boldsymbol{p}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})(\tilde{\boldsymbol{q}})$, which contradicts (4.42). Therefore, (4.44) holds. Thirdly, we shall show that
@@
$\langle\bm{\xi},M\bm{\xi}\rangle>0,\quad\forall\,\bm{\xi}\in{\mathcal{Q}}_{j_{\iota}}^{\circ}\setminus\{{\boldsymbol{0}}\}.$ (4.45)
@@
Since ${\mathcal{Q}}_{j_{\iota}}$ is a self-dual cone, one has ${\mathcal{Q}}_{j_{\iota}}^{\circ}=-{\mathcal{Q}}_{j_{\iota}}$. On the one hand, if there exists a nonzero vector $\hat{\bm{\xi}}\in-{\sf bdry}{\mathcal{Q}}_{j_{\iota}}$ such that $\langle\hat{\bm{\xi}},M\hat{\bm{\xi}}\rangle\leq 0$, then we know from Part (6d) of Lemma 3.2 that $-M\bm{\xi}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})(\bm{\xi})$, which contradicts (4.42). Therefore, it holds that
@@
$\langle\bm{\xi},M\bm{\xi}\rangle>0,\quad\forall\,\bm{\xi}\in{\sf bdry}{\mathcal{Q}}_{j_{\iota}}^{\circ}\setminus\{{\boldsymbol{0}}\}.$ (4.46)

### v3

@@ 4.2 Deriving the strong second-order sufficient condition
$\begin{array}[]{ll}H^{\top}\Theta HP_{k}\tilde{\bm{\eta}}+\sum\limits_{i=k+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}\tilde{\boldsymbol{p}}^{j_{i}}\in{\sf ker}(P_{k}^{\top})={\sf span}(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}),\end{array}$
which contradicts (61). Therefore, (60) holds, and it is easy to see that $T_{k}$ is nonsingular.
In the following, we separate the proof of (56) into three cases:
@@
Let $k:=\iota-1$ in (58). According to (12) and Part (6e) of Lemma 3.2, one can fix all $p^{j_{i}}=0$ and all ${\boldsymbol{p}}^{j_{i}}={\boldsymbol{0}}$ in (58) for all $i\neq\iota$. By doing this one can obtain that
(62) $T_{\iota-1}{\bm{\eta}}\neq-P_{\iota-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}p,\quad\forall\,p\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{\iota}}P_{\iota-1}{\bm{\eta}}),\quad\forall\,{\bm{\eta}}\in{\mathds{R}}^{\ell-\iota+1}\setminus\{{\boldsymbol{0}}\}.$
From (51) one has ${\mathcal{G}}_{j_{\iota}}=((\dot{\mu}^{*})^{j_{\iota}};-(\bar{\bm{\mu}}^{*})^{j_{\iota}})^{\top}{\mathcal{J}}g^{j_{\iota}}({\boldsymbol{x}}^{*})H$. Then, by defining the vector ${\bm{\eta}}:=T_{\iota-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{\iota-1}]^{\top}$ one can get
(63) $\langle(\dot{\mu}^{*})^{j_{\iota}};-(\bar{\bm{\mu}}^{*})^{j_{\iota}}),{\mathcal{J}}g^{j_{\iota}}({\boldsymbol{x}}^{*})HP_{{\iota}-1}{\bm{\eta}}\rangle={\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}{\bm{\eta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\leq 0.$
It can be easily observed from (12) that $p:=-1\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}{\bm{\eta}})$. Meanwhile, one has $T_{{\iota}-1}{\bm{\eta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}=-P_{{\iota}-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}p$. However, this contradicts (62), so that (63) fails. Therefore, (56) holds for all $i\in\{1,\ldots,|J_{4}|\}$.
@@
Let $k:=\iota-1$ in (58). According to (12) and Part (6e) of Lemma 3.2, one can fix all $p^{j_{i}}=0$ and all ${\boldsymbol{p}}^{j_{i}}={\boldsymbol{0}}$ in (58) for all $i\neq\iota$. By doing this one can obtain that
(64) $T_{\iota-1}{\bm{\eta}}\neq-P_{\iota-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}p,\quad\forall\,p\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{\iota}}P_{\iota-1}{\bm{\eta}}),\quad\forall\,{\bm{\eta}}\in{\mathds{R}}^{\ell-\iota+1}\setminus\{{\boldsymbol{0}}\}.$
From (51) one has ${\mathcal{G}}_{j_{\iota}}=(\dot{g}^{j_{\iota}}({\boldsymbol{x}}^{*});-\bar{g}^{j_{\iota}}({\boldsymbol{x}}^{*}))^{\top}{\mathcal{J}}g^{j_{\iota}}({\boldsymbol{x}}^{*})H$. Then, by letting ${\bm{\eta}}:=T_{\iota-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{\iota-1}]^{\top}$ one can get
(65) $(\dot{g}^{j_{\iota}}({\boldsymbol{x}}^{*});-\bar{g}^{j_{\iota}}({\boldsymbol{x}}^{*}))^{\top}{\mathcal{J}}g^{j_{\iota}}({\boldsymbol{x}}^{*})HP_{{\iota}-1}{\bm{\eta}}={\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}{\bm{\eta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\leq 0.$
Thus, it can be easily observed from (12) that $p:=-1\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}{\bm{\eta}})$. Moreover, by the definition of ${\bm{\eta}}$ we have $T_{{\iota}-1}{\bm{\eta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}=P_{{\iota}-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}=-P_{{\iota}-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}p$. However, this contradicts (64), so that (65) fails. Therefore, (56) holds for all $i\in\{|J_{4}|+1,\ldots,|J_{4,5}|\}$.
@@
(66) $[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\not\succ O.$
Let $k:=\iota-1$ in (60). According to Part (6e) of Lemma 3.2 (or (12) if $r_{j_{\iota}=0}$), one can fix ${\boldsymbol{p}}^{j_{i}}={\boldsymbol{0}}$ in (60) for all $i\neq\iota$. By doing this and using the fact that $\tilde{k}=k+\sum_{t=|J_{4,5}|+1}^{k}r_{j_{t}}$ one can obtain that
(67) $T_{\iota-1}{\bm{\eta}}\neq-P_{\iota-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}{\boldsymbol{p}},\quad\forall\,{\boldsymbol{p}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{\iota}}P_{\iota-1}{\bm{\eta}}),\quad\forall\,{\bm{\eta}}\in{\mathds{R}}^{\ell-\tilde{k}}\setminus\{{\boldsymbol{0}}\}.$
Note that ${\sf ker}P_{\iota-1}^{\top}={\sf span}\big{(}\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,\iota-1}\big{)}$. Moreover, recall from (53) and (54) that each $P_{i}$ has full column rank. Suppose that $[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}$ is singular, i.e., there exists a nonzero vector $\hat{\bm{\theta}}\in{\mathds{R}}^{1+r_{j_{\iota}}}$ such that $[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\hat{\bm{\theta}}={\boldsymbol{0}}.$ Then, on the one hand, if $\hat{\bm{\eta}}:=T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\hat{\bm{\theta}}={\boldsymbol{0}}$, one has $P^{\top}_{{\iota}-1}{\mathcal{G}}_{j_{\iota}}^{\top}\hat{\bm{\theta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\hat{\bm{\theta}}={\boldsymbol{0}}$, so that
@@
but this contradicts (46) of Corollary 4.11. On the other hand, if $\hat{\bm{\eta}}\neq{\boldsymbol{0}}$, one has from the fact ${\mathcal{G}}_{j_{\iota}}P_{\iota-1}\hat{\bm{\eta}}={\boldsymbol{0}}$ and Part (6a) of Lemma 3.2 (or (12) if $r_{j}=0$) that $-\hat{\bm{\theta}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{\iota}}P_{\iota-1}\hat{\bm{\eta}})$. Moreover, one can see that $T_{\iota-1}\hat{\bm{\eta}}=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}\hat{\bm{\theta}}=-P_{\iota-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}(-\hat{\bm{\theta}})$, which contradicts (67). Therefore, $[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}$ is nonsingular. Then, one can define the matrix $M:=([{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top})^{-1}$.
@@
For a vector ${\boldsymbol{q}}\in{\mathds{R}}^{1+r_{j_{\iota}}}$, one can define ${\boldsymbol{p}}:=-M{\boldsymbol{q}}$ and ${\bm{\eta}}:=-T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}{\boldsymbol{p}}$. Then one has ${\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}{\bm{\eta}}={\boldsymbol{q}}$ and, consequently, $T_{{\iota}-1}{\bm{\eta}}=-[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}{\boldsymbol{p}}=-P_{{\iota}-1}^{\top}{\mathcal{G}}_{j_{\iota}}^{\top}{\boldsymbol{p}}$. Moreover, since (67) holds, one obtains that
(68) $-M{\boldsymbol{q}}\notin{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})({\boldsymbol{q}}),\quad\forall{\boldsymbol{q}}\in{\mathds{R}}^{1+r_{{j}_{\iota}}}.$
@@
Next, we show that (66) can not be true. We first consider the case that $r_{j_{\iota}}=0$. If (66) holds in this case, by setting $q=[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}<0$, one has $-M^{-1}q=-1<0$. Then, by (12) we know that $-Mq\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}(0,0)(q)$, which contradicts (68). Consequently, one has $[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]T_{{\iota}-1}^{-1}[{\mathcal{G}}_{j_{\iota}}P_{{\iota}-1}]^{\top}>0$ in this case. In the following, we deal with the case that $r_{j_{\iota}}\geq 1$. Firstly, we shall prove that
@@
(69) $\langle\bm{\theta},M^{-1}\bm{\theta}\rangle>0,\quad\forall\,\bm{\theta}\in{\sf bdry}{\mathcal{Q}}_{{j_{\iota}}}\setminus\{{\boldsymbol{0}}\}.$
@@
Suppose on the contrary that there exists a vector $\hat{\bm{\theta}}\in{\sf bdry}{\mathcal{Q}}_{{j_{\iota}}}\setminus\{{\boldsymbol{0}}\}$ such that $\langle\hat{\bm{\theta}},M^{-1}\hat{\bm{\theta}}\rangle\leq 0.$ By taking $\hat{\boldsymbol{q}}:=M^{-1}\bf\hat{\bm{\theta}}$, one has $\langle-\hat{\bm{\theta}},\hat{\boldsymbol{q}}\rangle\geq 0$. Then, by Part (6c) of Lemma 3.2 we know that $-M\hat{\boldsymbol{q}}=-\hat{\bm{\theta}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})(\hat{\boldsymbol{q}})$, which contradicts (68). Therefore, (69) holds. Secondly, we shall show that the following holds:
(70) $\langle\bm{\theta},M^{-1}\bm{\theta}\rangle>0,\quad\forall\,\bm{\theta}\in{\sf int}{\mathcal{Q}}_{j_{\iota}}.$
@@
For the purpose of contradiction, assume that (70) does not hold. Since the matrix $M^{-1}$ is nonsingular and symmetric, we know from (69) and Lemma 2.1 that there exists a nonzero vector $\tilde{\boldsymbol{p}}\in-{\sf int}{\mathcal{Q}}_{j_{\iota}}$ such that $\tilde{\boldsymbol{q}}:=-M^{-1}\tilde{\boldsymbol{p}}\in-{\mathcal{Q}}_{j_{\iota}}$. From Part (6b) of Lemma 3.2 one can see that $-M\tilde{\boldsymbol{q}}=\tilde{\boldsymbol{p}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})(\tilde{\boldsymbol{q}})$, which contradicts (68). Therefore, (70) holds. Thirdly, we shall show that
@@
(71) $\langle\bm{\xi},M\bm{\xi}\rangle>0,\quad\forall\,\bm{\xi}\in{\mathcal{Q}}_{j_{\iota}}^{\circ}\setminus\{{\boldsymbol{0}}\}.$
@@
Since ${\mathcal{Q}}_{j_{\iota}}$ is a self-dual cone, one has ${\mathcal{Q}}_{j_{\iota}}^{\circ}=-{\mathcal{Q}}_{j_{\iota}}$. On the one hand, if there exists a nonzero vector $\hat{\bm{\xi}}\in-{\sf bdry}{\mathcal{Q}}_{j_{\iota}}$ such that $\langle\hat{\bm{\xi}},M\hat{\bm{\xi}}\rangle\leq 0$, then we know from Part (6d) of Lemma 3.2 that $-M\bm{\xi}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{\iota}}}({\boldsymbol{0}},{\boldsymbol{0}})(\bm{\xi})$, which contradicts (68). Therefore, it holds that
@@
(72) $\langle\bm{\xi},M\bm{\xi}\rangle>0,\quad\forall\,\bm{\xi}\in{\sf bdry}{\mathcal{Q}}_{j_{\iota}}^{\circ}\setminus\{{\boldsymbol{0}}\}.$